OpenStudy (anonymous):

if a=-1 and b=5 what is 5a-7b+b^2 isnt it 55 and also i can not figure out 9xy^2-11xy^2

OpenStudy (cwrw238):

5* -1 - 7* 5 + 5^2 = ?

OpenStudy (anonymous):

55

OpenStudy (cwrw238):

no take the first 2 terms -1* 5 - 7*5 = -5 - 35 = ?

sammixboo (sammixboo):

Just plug in the values and follow the Order of Operations P - Do what is in the PARENTHESIS first E - Now do the exponent(s) M/D - Do multiplication or division in the order starting from the LEFT side of the equation and end at the RIGHT of the equation A/S Do addition or subtraction in the order starting from the LEFT side of the equation and end at the RIGHT of the equation

OpenStudy (anonymous):

-40

OpenStudy (cwrw238):



OpenStudy (cwrw238):

right now its -40 + 5^2 now as sammixboo indicated do the exponent part first ( the 5^2) then add

OpenStudy (cwrw238):

ok?

OpenStudy (cwrw238):

what is 5^2?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

-15 is the answer i got....

OpenStudy (cwrw238):

ok now we have -40 + 25 = ?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

now 9xy^2-11xy^2

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

i guess you want to simplify this notice that these are 2 like terms they both have xy^2 in them 9 - 11 = -2 so 9xy^2 - 11xy^2 = ?